OpenStudy (anonymous):

If ƒ(x ) = x 2 + 1 and g(x ) = 3x + 1, find [ƒ(2) - g(1)] 2.

OpenStudy (anonymous):

f(x)=x^2+1 f(2)=2^2+1=5 put in 2 where you see x g(x)=3x+1 g(1)=3+1=4 f(2)-g(1)=5-4=1 the 2 outside means multiply by 2 =1*2=2